OpenStudy (anonymous):

Mario’s company makes unusually shaped imitation gemstones. One gemstone had 12 faces and 10 vertices. How many edges did the gemstone have? A. 23 edges B. 22 edges C. 25 edges D. 20 edges

OpenStudy (anonymous):

from euler's formula faces+vertices=edges+2, this gives 12+10=e+2 then e=20.

OpenStudy (anonymous):

thank you :)